OpenStudy (anonymous):

Please help...I am so confused....(a^1/2 b^-1/3)(ab)

OpenStudy (anonymous):

(a^1/2)(a) = a^(3/2) (b^(-1/3))(b) = b^(2/3) Final answer: a^(3/2)b^(2/3) Keep in mind that a^x(a^y) = a^(x+y_

OpenStudy (anonymous):

thank you I'm still confused on what you did to get the answer you got..Sorry i'm not very good at this

OpenStudy (anonymous):

I have used the law of exponents which says that a^x times a^y = a^ (x+y) A base to an exponent multiplied by the same base to another exponent, then you get the base raised to the sum of the two exponents. x^2 times x^3 = x^5.

OpenStudy (anonymous):

so a^1/2 times a = a^(3/2) because you add the exponents 1/2 and 1 = 3/2

OpenStudy (anonymous):

a is like a^1